# CorDA: Context-Oriented Decomposition Adaptation of Large Language Models

### v1

@@ 4 Experiments
In experiments, we finetune the pre-trained large language model LLaMA-2-7b [54] on Math, Code, and Instruction Following tasks, and also apply our method to the General Language Understanding Evaluation (GLUE) benchmark with RoBERTabase [39]. The world knowledge is evaluated by the exact match scores (%) on TriviaQA [23], NQ open [29], and WebQS [4]. Following the settings in [43], the Math ability is trained on MetaMathQA [66] and tested on GSM8k [9] and Math [66] validation sets. Code is trained on CodeFeedback [77] and tested on HumanEval [8] and MBPP [3]. Instruction following is trained on WizardLM-Evol-Instruct [59] and tested on MTBench [76]. The complete implementation details are described in Appendix A.
@@ 4.2 Knowledge-Preserved Adaptation Results
Table 1: The experimental results of CorDA in the knowledge-preserved adaptation mode and comparison with full finetuning, LoRA, and PiSSA. LLaMA-2-7B is used to finetune on (a) Math, (b) Code, and (c) Instruction Following tasks. The rank $r$ of LoRA, PiSSA, and CorDA is 128. CorDA is initialized with the NQ open samples to collect the covariance matrices. All methods are implemented by us under the same training and evaluation settings. The row of “LLaMA-2-7B” shows the world knowledge performance of the original pre-trained model.
@@
We finetune LLaMA-2-7B with full finetuning, LoRA, PiSSA, and our proposed CorDA on Math, Code, and Instruction Following tasks. In the knowledge-preserved adaptation mode, we randomly sample 256 questions from the NQ open training set and collect covariance matrices to initialize the adapters by Eq. (4). We report the finetuning performance and also the world knowledge performance of the finetuned model to manifest the overall ability of both new task learning and world knowledge maintaining. As shown in Table 1, after finetuning on Math, all the three compared methods, full finetuning, LoRA, and PiSSA, suffer from drastic performance drop on TriviaQA and NQ open. Especially on NQ open, the ability is almost lost. PiSSA achieves the best accuracies on both GSM8k and Math, but meanwhile has the lowest results on the world knowledge benchmarks among the four methods. As a comparison, our method not only enjoys better finetuning performances than LoRA, but also achieves the best results on the three world knowledge benchmarks.
A similar pattern can be also observed in the finetuning of Code. As shown in Table 1, full finetuning has the best ability on HumanEval and MBPP, but is the lowest on world knowledge performance. It is understandable that CorDA in knowledge-preserved adaptation is not as advantageous as PiSSA for finetuning performance, because PiSSA uses the largest singular values and their singular vectors as the adapter that dominates the weight update, while we keep them frozen and adapt the smallest components. Nevertheless, our method has the best average score in all the three finetuning tasks. A surprising result is achieved by our method in instruction following in Table 1, where CorDA surpasses all the three compared methods in both world knowledge performance and the new ability evaluated by MTBench. The world knowledge on NQ open is almost intact (14.43%) compared to the original performance (14.99%) before finetuning. These results reveal that our knowledge-preserved adaptation is an effective way to mitigate world knowledge forgetting and improve the overall ability.
@@ 4.3 Instruction-Previewed Adaptation Results
When the goal is to learn the target task as much as possible without concerning the loss of world knowledge, CorDA in the instruction-previewed adaptation mode satisfies this demand as it is able to further strengthen the finetuning performance. In this mode, we randomly sample instruction and response from the training data to collect the covariance matrices, and adopt the largest $r$ components to initialize the adapters by Eq. (5). We compare our method with full finetuning, LoRA, DoRA, and PiSSA in the same three tasks, Math, Code, and Instruction Following. Compared with the knowledge-preserved adaptation in Table 1, CorDA in the instruction-previewed adaptation mode largely improves the finetuning performance on the five benchmarks, as shown in Table 2. Concretely, our method achieves the best performance on GSM8k, Math, MTBench, and the average score. Compared with PiSSA that adopts a similar adapter design but with no data context, our method has better results on all the five benchmarks. The training loss curves on Math are shown in Figure 3, where CorDA converges at a lower loss than LoRA and PiSSA in both $r=128$ and $r=32$. Only CorDA exhibits an obvious lower loss than full finetuning when $r=128$. These results corroborate the benefits of data context to the initialized adapter, i.e., the pre-captured task characteristic is able to accommodate the new ability and lead to a better performance.
@@
Figure 3: The training loss curves on MetaMath of full finetuning, LoRA, PiSSA, and CorDA with (a) rank 128 and (b) rank 32. The corresponding accuracies on GSM8k and Math are reported on the legends. Smoothing is performed for the loss curves.
Table 2: The experimental results of CorDA in the instruction-previewed adaptation mode on Math, Code, and Instruction Following tasks using LLaMA-2-7B. CorDA is initialized with samples from each of the finetuning datasets (MetaMathQA, CodeFeedback, and WizardLM-Evol-Instruct) for the three tasks, respectively. The rank $r$ of LoRA, DoRA, PiSSA, and CorDA is 128. All methods are implemented by us under the same training and evaluation settings.
@@
Table 3: The experimental results of CorDA in the instruction-previewed adaptation mode on the GLUE benchmark using RoBERTabase. CorDA is initialized with samples from each of the finetuning datasets. The rank $r$ of LoRA, DoRA, and CorDA is 128. All methods are implemented by us under the same training and evaluation settings. Matthew’s correlation and Pearson’s correlation are the metrics of CoLA and STS-B, respectively. The metric of the other tasks is accuracy.
@@
We also apply our method to the General Language Understanding Evaluation (GLUE) benchmark [56] by finetuning the RoBERTabase model [39]. We adopt LoRA, DoRA, and our method with a rank of $128$ for all linear layers in the model except the classification head. For our method, we sample train data from each of the tasks to initialize adapters in the instruction-previewed mode and finetune the corresponding task. As shown in Table 3, our method achieves the best performance on the MRPC, CoLA, and RTE tasks, and the highest average score.
@@ 4.4 Discussions
Limitations. The two adaptation modes developed in this paper highlight different aspects in usage. However, the knowledge-preserved mode, while being adept at maintaining world knowledge, is naturally not advantageous on finetuning performance compared with the instruction-previewed mode. How to develop an initialization strategy [63] for adapters combining the merits of the two modes to maximize both objectives deserves future exploration.
@@ 5 Conclusion
In this paper, we propose a new parameter-efficient fine-tuning method, named context-oriented decomposition adaptation (CorDA). It performs singular value decomposition for pre-trained weights oriented by the covariance matrix that captures the context of the task concerned, and aggregates the context into the principle components for maintaining or adapting. Accordingly, our method is able to support two implementation modes, the knowledge-preserved adaptation to mitigate world knowledge forgetting and the instruction-previewed adaptation for better finetuning performance. In experiments, our knowledge-preserved adaptation not only achieves better finetuning performance than LoRA, but also maintains the world knowledge well, leading to the best average scores on three finetuning tasks. Our instruction-previewed adaptation is able to further enhance the finetuning performance, surpassing the state-of-the-art parameter-efficient fine-tuning methods DoRA and PiSSA.
@@ A.1 Fientuning on Math, Code, and Instruction Following
For finetuning tasks on Math, Code, and Instruction Following, we adopt the same training setting as PiSSA [43]. Concretely, optimization is performed with the AdamW optimizer, a batch size of 128, and a learning rate of 2e-5. We employ cosine annealing schedules with a warmup ratio of 0.03 and do not apply weight decay. Training is conducted exclusively on the first 100,000 conversations from the dataset for one epoch, with loss computation solely based on the response. Our experiments are executed on the NVIDIA A100-SXM4(40/80GB) GPUs. Publicly available platforms are utilized for the evaluation of world knowledge (TriviaQA, NQ open, and Web QS)  [^2] https://github.com/EleutherAI/lm-evaluation-harness, Code (HumanEval and MBPP)  [^3] https://github.com/bigcode-project/bigcode-evaluation-harness, and Instruction Following (MTBench)  [^4] https://github.com/lm-sys/FastChat.
@@ A.2 GLUE Benchmark
To ensure fair comparison across Full Finetuning, LoRA, DoRA, and CorDA in the GLUE benchmark, we implement all methods under the same training and evaluation settings. The AdamW optimizer is used with a batch size of 32 and a learning rate of 4e-5 for 3 epochs, following a linear learning rate schedule. The max token length is set as 128. The rank of LoRA, DoRA, and our CorDA is 128. For covariance matrix collection of CorDA, we concatenate the representative content of each training sample to form a text sequence. From this sequence, 256 text segments, each containing 256 tokens, are randomly sampled. The selected content for each task is as follows: MRPC, RTE, and STS-B using “sentence1”, CoLA and SST-2 using “sentence”, and QNLI using “question”. All methods are trained on a single NVIDIA A100-SXM4(40/80GB) GPU.

### v2

@@ 4 Experiments
In experiments, we fine-tune the pre-trained large language model LLaMA-2-7b [58] on Math, Code, and Instruction Following tasks, and also apply our method to the General Language Understanding Evaluation (GLUE) benchmark with RoBERTabase [43]. The world knowledge is evaluated by the exact match scores (%) on TriviaQA [23], NQ open [29], and WebQS [4]. Following the settings in [47], the Math ability is trained on MetaMathQA [71] and tested on GSM8k [9] and Math [71] validation sets. Code is trained on CodeFeedback [82] and tested on HumanEval [8] and MBPP [3]. Instruction following is trained on WizardLM-Evol-Instruct [63] and tested on MTBench [81]. The complete implementation details are described in Appendix A.
(a) Perplexity on Wikitext-2
(b) Perplexity on PTB
@@ 4.2 Knowledge-Preserved Adaptation Results
Table 1: The experimental results of CorDA in the knowledge-preserved adaptation mode and comparison with full fine-tuning, LoRA, and PiSSA. LLaMA-2-7B is used to fine-tune on (a) Math, (b) Code, and (c) Instruction Following tasks. The rank $r$ of LoRA, PiSSA, and CorDA is 128. CorDA is initialized with the NQ open samples to collect the covariance matrices. All methods are implemented by us under the same training and evaluation settings. The row of “LLaMA-2-7B” shows the world knowledge performance of the original pre-trained model.
@@
We fine-tune LLaMA-2-7B with full fine-tuning, LoRA, PiSSA, and our proposed CorDA on Math, Code, and Instruction Following tasks. In the knowledge-preserved adaptation mode, we randomly sample 256 questions from the NQ open training set and collect covariance matrices to initialize the adapters by Eq. (4). We report the fine-tuning performance and also the world knowledge performance of the fine-tuned model to manifest the overall ability of both new task learning and world knowledge maintaining. As shown in Table 1, after fine-tuning on Math, all the three compared methods, full fine-tuning, LoRA, and PiSSA, suffer from drastic performance drop on TriviaQA and NQ open. Especially on NQ open, the ability is almost lost. PiSSA achieves the best accuracies on both GSM8k and Math, but meanwhile has the lowest results on the world knowledge benchmarks among the four methods. As a comparison, our method not only enjoys better fine-tuning performances than LoRA, but also achieves the best results on the three world knowledge benchmarks.
A similar pattern can be also observed in the fine-tuning of Code. As shown in Table 1, full fine-tuning has the best ability on HumanEval and MBPP, but is the lowest on world knowledge performance. It is understandable that CorDA in knowledge-preserved adaptation is not as advantageous as PiSSA for fine-tuning performance, because PiSSA uses the largest singular values and their singular vectors as the adapter that dominates the weight update, while we keep them frozen and adapt the smallest components. Nevertheless, our method has the best average score in all the three fine-tuning tasks. A surprising result is achieved by our method in instruction following in Table 1, where CorDA surpasses all the three compared methods in both world knowledge performance and the new ability evaluated by MTBench. The world knowledge on NQ open is almost intact (14.43%) compared to the original performance (14.99%) before fine-tuning. These results reveal that our knowledge-preserved adaptation is an effective way to mitigate world knowledge forgetting and improve the overall ability.
@@ 4.3 Instruction-Previewed Adaptation Results
When the goal is to learn the target task as much as possible without concerning the loss of world knowledge, CorDA in the instruction-previewed adaptation mode satisfies this demand as it is able to further strengthen the fine-tuning performance. In this mode, we randomly sample instruction and response from the training data to collect the covariance matrices, and adopt the largest $r$ components to initialize the adapters by Eq. (5). We compare our method with full fine-tuning, LoRA, DoRA, and PiSSA in the same three tasks, Math, Code, and Instruction Following. Compared with the knowledge-preserved adaptation in Table 1, CorDA in the instruction-previewed adaptation mode largely improves the fine-tuning performance on the five benchmarks, as shown in Table 2. Concretely, our method achieves the best performance on GSM8k, Math, MTBench, and the average score. Compared with PiSSA that adopts a similar adapter design but with no data context, our method has better results on all the five benchmarks. The training loss curves on Math are shown in Figure 3, where CorDA converges at a lower loss than LoRA and PiSSA in both $r=128$ and $r=32$. Only CorDA exhibits an obvious lower loss than full fine-tuning when $r=128$. These results corroborate the benefits of data context to the initialized adapter, i.e., the pre-captured task characteristic is able to accommodate the new ability and lead to a better performance.
@@
Figure 3: The training loss curves on MetaMath of full fine-tuning, LoRA, PiSSA, and CorDA with (a) rank 128 and (b) rank 32. The corresponding accuracies on GSM8k and Math are reported on the legends. Smoothing is performed for the loss curves.
Table 2: The experimental results of CorDA in the instruction-previewed adaptation mode on Math, Code, and Instruction Following tasks using LLaMA-2-7B. CorDA is initialized with samples from each of the fine-tuning datasets (MetaMathQA, CodeFeedback, and WizardLM-Evol-Instruct) for the three tasks, respectively. The rank $r$ of LoRA, DoRA, PiSSA, and CorDA is 128. All methods are implemented by us under the same training and evaluation settings.
@@
Table 3: The experimental results of CorDA in the instruction-previewed adaptation mode on the GLUE benchmark using RoBERTabase. CorDA is initialized with samples from each of the fine-tuning datasets. The rank $r$ of LoRA, DoRA, and CorDA is 128. All methods are implemented by us under the same training and evaluation settings. Matthew’s correlation and Pearson’s correlation are the metrics of CoLA and STS-B, respectively. The metric of the other tasks is accuracy.
@@
We also apply our method to the General Language Understanding Evaluation (GLUE) benchmark [60] by fine-tuning the RoBERTabase model [43]. We adopt LoRA, DoRA, and our method with a rank of $128$ for all linear layers in the model except the classification head. For our method, we sample train data from each of the tasks to initialize adapters in the instruction-previewed mode and fine-tune the corresponding task. As shown in Table 3, our method achieves the best performance on the MRPC, CoLA, and RTE tasks, and the highest average score.
@@ 4.4 Discussions
Limitations. The two adaptation modes developed in this paper highlight different aspects in usage. However, the knowledge-preserved mode, while being adept at maintaining world knowledge, is naturally not advantageous on fine-tuning performance compared with the instruction-previewed mode. How to develop an initialization strategy [68] for adapters combining the merits of the two modes to maximize both objectives deserves future exploration.
@@ 5 Conclusion
In this paper, we propose a new parameter-efficient fine-tuning method, named context-oriented decomposition adaptation (CorDA). It performs singular value decomposition for pre-trained weights oriented by the covariance matrix that captures the context of the task concerned, and aggregates the context into the principle components for maintaining or adapting. Accordingly, our method is able to support two implementation modes, the knowledge-preserved adaptation to mitigate world knowledge forgetting and the instruction-previewed adaptation for better fine-tuning performance. In experiments, our knowledge-preserved adaptation not only achieves better fine-tuning performance than LoRA, but also maintains the world knowledge well, leading to the best average scores on three fine-tuning tasks. Our instruction-previewed adaptation is able to further enhance the fine-tuning performance, surpassing the state-of-the-art parameter-efficient fine-tuning methods DoRA and PiSSA.
Acknowledgments and Disclosure of Funding
The research reported in this publication was supported by funding from King Abdullah University of Science and Technology (KAUST) - Center of Excellence for Generative AI, under award number 5940.
This work was also supported in part by the National Natural Science Foundation of China under Grant 62376069, in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, and in part by Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012027.
@@ A.1 Fientuning on Math, Code, and Instruction Following
For fine-tuning tasks on Math, Code, and Instruction Following, we adopt the same training setting as PiSSA [47]. Concretely, optimization is performed with the AdamW optimizer, a batch size of 128, and a learning rate of 2e-5. We employ cosine annealing schedules with a warmup ratio of 0.03 and do not apply weight decay. Training is conducted exclusively on the first 100,000 conversations from the dataset for one epoch, with loss computation solely based on the response. Our experiments are executed on the NVIDIA A100-SXM4(40/80GB) GPUs. Publicly available platforms are utilized for the evaluation of world knowledge (TriviaQA, NQ open, and Web QS)  [^2] https://github.com/EleutherAI/lm-evaluation-harness, Code (HumanEval and MBPP)  [^3] https://github.com/bigcode-project/bigcode-evaluation-harness, and Instruction Following (MTBench)  [^4] https://github.com/lm-sys/FastChat.
@@ A.2 GLUE Benchmark
To ensure fair comparison across Full fine-tuning, LoRA, DoRA, and CorDA in the GLUE benchmark, we implement all methods under the same training and evaluation settings. The AdamW optimizer is used with a batch size of 32 and a learning rate of 4e-5 for 3 epochs, following a linear learning rate schedule. The max token length is set as 128. The rank of LoRA, DoRA, and our CorDA is 128. For covariance matrix collection of CorDA, we concatenate the representative content of each training sample to form a text sequence. From this sequence, 256 text segments, each containing 256 tokens, are randomly sampled. The selected content for each task is as follows: MRPC, RTE, and STS-B using “sentence1”, CoLA and SST-2 using “sentence”, and QNLI using “question”. All methods are trained on a single NVIDIA A100-SXM4(40/80GB) GPU.
@@ Appendix B Appendix: More Results
Figure 4: Covariance visualization results for “self_attn.k_proj”, “self_attn.o_proj”, “mlp.down_proj”, and “mlp.gate_proj” weights in the 0-th layer. Please zoom in for a better view.
Figure 5: Covariance visualization results for “self_attn.o_proj” weights in different depth layers. Please zoom in for a better view.
Appendix C Appendix: Covariance Visualization Results
We provide the visualization results of the covariance matrices collected from three tasks MetaMath, NQ open, and Trivia QA in Figures 4 and 5.
Since the original dimension in 4096 or 11008 will be too large to be informative, we downsample the covariance matrices into $32\times 32$ and visualize their heatmaps. We provide the results from the activations before different linear weights including self_attn.k_proj (the same as self_attn.q_proj and self_attn.v_proj due to the same input), self_attn.o_proj, mlp.down_proj, and mlp.gate_proj (the same as mlp.up_proj) in the first layer, and the self_attn.o_proj weight in later layers. It is shown that the heatmaps from NQopen and TriviaQA (both are QA tasks) share some similar patterns (marked in red circles), which do not appear in the heatmap from the different task MetaMath. Therefore, when the inputs of different tasks are fed into an LLM, the covariance matrix from activations will exhibit different patterns. The visualization result empirically supports that the covariance matrix patterns can be used to characterize the triggered task. We use such patterns to orientate the decomposition of LLM pretrained weights, to make the resulting adapter initialization task-dependent.